# MIM-Refiner: A Contrastive Learning Boost from Intermediate Pre-Trained Representations

### v2

@@ 6 Conclusion
We introduce MIM-Refiner, a procedure to refine pre-trained MIM models. Motivated by the insights that the representation quality of MIM models peaks in the middle of the encoder, we employ an ensemble of instance discrimination heads attached at multiple blocks towards the end of the encoder, including the blocks where representation quality peaks, to improve upon the best existing representation. We train this ensemble for a short duration to improve the representation for downstream tasks such as classification, clustering or semantic segmentation. As instance discrimination objective we introduce Nearest Neighbor Alignment, an adaption of existing nearest neighbor based objectives that uses the nearest neighbor relation only for aligning positive pairs and not for pushing negative pairs apart, to improve the signal quality from negatives.
@@
7 Acknowledgements
@@
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, Leonardo at CINECA, Italy, MeluXina at LuxProvide, Luxembourg and LUMI at CSC, Finland.
@@
Appendix A Reproducibility
We publish the full codebase used for this work which contains code for training the models, evaluations and dataset preparation. https://github.com/ml-jku/MIM-Refiner
Appendix B Limitations
One limitation of MIM-Refiner is that it requires batch normalization [39] layers in the ID head. Without them, performance decreases significantly. As we do not change anything in the MIM pre-training, we hypothesize that feature statistics after MIM pre-training are not suited for an ID task and therefore require per-feature normalization. The batch normalization layers significantly decrease scalability to distributed hardware setups as each layer requires a synchronization of batch statistics.
MIM-Refiner addresses a common issue with MIM models: their typically lightweight decoder often delegates part of the reconstruction to the encoder, resulting in subpar representations for the later encoder blocks in downstream tasks. Alternatively, one could simply argue for a larger decoder. However, a larger decoder increases computational costs since the decoder typically operates on the full sequence length. Additionally, the direction of a larger decoder was explored to a certain extent in the original MAE paper [29], where larger decoders performed worse in fine-tuning and linear probing. Successor models such as CrossMAE perform well with a deeper decoder but still show decreasing representation quality in later layers, as shown in Appendix C.1.
@@ Appendix B Limitations
Appendix C Extended Results
C.1 Intermediate Representation Analysis of Additional MIM Methods
(a) MAE [29]
(b) data2vec 2.0 [6]
(c) CrossMAE [27]
(d) dBOT [50]
(e) CAE [15]
Figure 6: Intermediate representation analysis of various MIM models. As models get bigger, the $k$-NN accuracy degrades more towards the end of the encoder, especially for L/16 and H/14 models. This pattern is consistent over various MIM models.
C.2 Where to Attach ID Heads?
Table 6(b) ablates different choices of where to attach ID heads on a D2V2 pre-trained ViT-L/16 and ViT-H/14. On ViT-L/16, there is almost no difference of where to attach the ID heads in the last third of blocks. We tried to transfer this insight to ViT-H/14 where the default setting of attaching ID heads to all later blocks performs better. We therefore use the default setting of attaching ID heads to the last 8 blocks (ViT-L and ViT-2B) or the last 12 blocks (ViT-H).
@@ C.2 Where to Attach ID Heads?
C.3 Low-shot Classification Without Color Augmentations
@@
Table 7: ImageNet-1K low-shot and feature evaluations of a refined data2vec 2.0 ViT-H/14 with and without color augmentations. Omitting color augmentations (i.e. using only crop & flip augmentations) improves ImageNet-1K low-shot performance.
@@
C.4 Extended ImageNet-1K Low-shot and Feature Evaluations
Table 8 extends Table 2 with additonal results and comparison to more SSL models.
@@ C.4 Extended ImageNet-1K Low-shot and Feature Evaluations
L/16 CrossMAE [27] 16.8 34.0 52.4 63.2 77.7 74.4 53.4
MAE [29] 14.3 34.9 56.9 67.7 79.3 77.5 60.6
dBOT [50] 28.0 46.9 62.4 70.0 80.2 77.8 61.3
D2V2 [6] 24.1 58.8 72.1 75.1 81.5 78.2 51.8
CAE [15] 28.1 56.5 68.4 71.4 79.5 80.0 66.9
iBOT [81] 48.5 58.2 66.5 73.3 79.0 81.1 78.0
Mugs [82] 52.9 62.3 69.4 76.2 80.3 82.1 80.4
@@
g/14 DINOv2 [56] (142M) 60.5 68.3 74.4 79.1 83.8 86.5 83.5
C.5 Extended ImageNet-1K Cluster Evaluations
@@
Figure 7: Block-wise cluster analysis for refined and unrefined MAE and D2V2 (H/14). Upper Left: Cluster accuracy w.r.t. lowest $k$-means loss per block. Upper Right: Silhouette score w.r.t. ground truth ImageNet-1K classes. Bottom: Pairwise cluster label similarity between subsequent blocks (details in Section C.5).
C.6 Extended Transfer Learning Results
@@
L/16 CrossMAE [27] 5.4 45.6 63.7 39.8 81.2 30.9
MAE [29] 7.1 51.6 68.9 42.8 83.0 33.6
dBOT [50] 7.1 53.1 71.2 44.6 83.4 34.1
D2V2 [6] 5.5 53.1 70.6 39.4 81.5 38.8
CAE [15] 7.4 54.5 71.2 46.4 84.8 36.5
iBOT [81] 15.8 51.5 65.5 56.0 87.6 35.6
Mugs [82] 19.5 53.2 66.9 61.5 87.9 34.8
@@
L/7 MSN [2] 17.0 38.0 48.1 - - -
C.7 VTAB Individual Dataset Results
@@
C.8 Extended Fine-tuning Results
Table 14 extends Table 5 with additional MIM models. MIM-Refiner consistently improves also on fine-tuning with an abundance of labels. Table 15 shows individual performances on robustness datasets for fine-tuned models on ImageNet-1K.
@@ C.8 Extended Fine-tuning Results
C.9 Fine-tuning ViT-2B Models
@@
C.10 Impact of Multi-Crop Augmentation
@@ C.11 High-dimensional $k$-NN
Appendix D Implementation Details
D.1 Evaluations
@@ D.4 ID Head Architecture
D.5 Refinement Hyperparameters
Hyperparameters for the refinement stage are listed in Table 18. Following MAE-CT [47], we initialize all ID heads first by training them with a frozen encoder to ensure a good learning signal from the start of the refinement process. For this initialization, we use the same hyperparameters as in Table 18 except that we use 20 epochs for all models, a learning rate of 2e-4 and a top1-NN lookup. As we do not use a momentum encoder during training, we instead track an EMA of the encoder and use the EMA then for downstream tasks. As ViT-2B is very expensive to train, we freeze the first 6 blocks (for refinement and also for evaluation). As shown in Table 4(b) this slightly reduces performance but also reduces memory consumption and runtime.
Table 18: MIM-Refiner hyperparameters.
@@ D.5 Refinement Hyperparameters
   Freeze Blocks 0 (everything exept ViT-2B), 6 (ViT-2B)
@@
   Local View Resolution 96 (L), 98 (H)
@@
Appendix E Evaluation Details
E.1 GPU Hours Benchmark
For benchmarking GPU hours, we follow the setup from MAE-CT [47]: we conduct a comparison by implementing all methods in a single code-base and conducting short training runs on a single A100 40GB PCIe card. These runs are executed in mixed precision training mode and with the highest possible batchsize that is a power of 2. The runtime of these benchmark runs is then extrapolated to the reported number of epochs. Benchmarks are conducted in pytorch 2.1 with CUDA 12.1. FLOPS are measured with the fvcore library [^1] https://github.com/facebookresearch/fvcore. For the 1-shot classification plot in Figure 2, we do not take into account that some models train on higher resolutions (e.g. DINOv2) for visual clarity.
E.2 MAE Intermediate Representation Analysis
@@
For the visualization in Figure 3(d), we calculate the delta from one block to the next and normalize the deltas by dividing by the maximum delta. We do this for both the $k$-NN accuracy and the reconstruction loss. Figure 8 shows the reconstruction loss per block and the same plot for a MAE L/16, where a similar behavior can be observed.
Figure 8: Left: Reconstruction loss per block of MAEs. Right Relative improvement of reconstruction loss and $k$-NN accuracy for a MAE L/16.
E.3 ImageNet-1K Low-shot Evaluation Details

### v4

@@ 6 Conclusion
We introduce MIM-Refiner, a procedure to refine pre-trained MIM models. Motivated by the insights that the representation quality of MIM models peaks in the middle of the encoder, we employ an ensemble of instance discrimination heads attached at multiple blocks towards the end of the encoder, including the blocks where representation quality peaks, to improve upon the best existing representation. We train this ensemble for a short duration to improve the representation for downstream tasks such as classification, clustering or semantic segmentation.
@@
Extensive evaluations show the potential of MIM-Refiner for training large-scale vision foundation models that provide general-purpose off-the-shelf features for a broad range of downstream tasks.
Acknowledgements
@@
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, LUMI at CSC, Finland and Leonardo at CINECA, Italy. We acknowledge access to LEONARDO at CINECA, Italy, via an AURELEO (Austrian Users at LEONARDO supercomputer) project.
@@
Ethics Statement
Our work proposes a training methodology that can readily reuse existing models, which makes use of the expended energy for large-scale pre-training of MIM models instead of training models from scratch, which would need much more energy than refining existing models. Additionally, our models learn strong general-purpose features that can be readily used for a broad range of downstream tasks without needing to fine-tune or retrain a model. This further reduces the energy consumption and carbon emissions required for training state-of-the-art models in computer vision tasks.
The strong general-purpose features of our models provide a fundamental advancement to computer vision, inheriting their potential benefits. On the one hand, our models could be used for environmental monitoring of endangered species where little data is available. On the other hand, the strong few-shot performance of our models could lead to overconfidence in rare disease classifiers where sufficient data is indispensable and checking with a healthcare professional is necessary to avoid misdiagnoses.
Appendix A Limitations
One limitation of MIM-Refiner is that it requires batch normalization [43] layers in the ID head. Without them, performance decreases significantly. Similar observations have been made in other contrastive learning literature [16]. The batch normalization layers significantly decrease scalability to distributed hardware setups as each layer requires a synchronization of batch statistics [14].
MIM-Refiner addresses a common issue with MIM models: their typically lightweight decoder often delegates part of the reconstruction to the encoder, resulting in subpar representations for the later encoder blocks in downstream tasks. Alternatively, one could simply argue for a larger decoder. However, a larger decoder increases computational costs since the decoder typically operates on the full sequence length. Additionally, the direction of a larger decoder was explored to a certain extent in the original MAE paper [31], where larger decoders performed worse in fine-tuning and linear probing. Successor models such as CrossMAE perform well with a deeper decoder but still show decreasing representation quality in later layers, as shown in Appendix B.4.
@@ Appendix A Limitations
To address the limitation of MIM-Refiner requiring batch normalization layers, we explore another variant that copies the weights of the intermediate block with the highest $k$-NN accuracy to all subsequent blocks and sets the weights/biases of the last projection in the copied attention/MLP blocks to 0. This is similar to the above mentioned approach (ii), except that the weights of the copied blocks are not random but copied from an intermediate block. We then train the model with only a single head attached at the last block. This drastically reduces the number of batch normalization layers. Table 6 shows that such an approach can yield competitive performances on smaller models, but is outperformed by MIM-Refiner on larger ones.
$k$-NN MAE L/16 MAE H/14
Copy Blocks 81.5 81.8
MIM-Refiner 81.5 82.3
Table 6: Copying the peak-representation block to subsequent blocks while setting the last attention/MLP projection weights/biases to 0 can improve scalability to even larger distributed setups due to requiring less batch normalization layers. This approach (“Copy Blocks”) is competitive to an ensemble of ID heads (“MIM-Refiner”) on smaller models but is worse on larger models.
Another approach to improve scalability to larger distributed systems is to only aggregate batch statistics within a node. This avoids costly inter-node communication and instead only requires intra-node connections which is typically much faster.
As MIM-Refiner is quite computationally efficient even with the batchnorm synchronization limitation, we do not explore these approaches further.
Additionally, MIM-Refiner explores efficiently scaling self-supervised models while keeping the pre-training dataset fixed to ImageNet-1K. ImageNet-1K is a relatively high-quality dataset with a class-balanced label distribution and commonly contains a single object located near the center of the image. Scaling MIM-Refiner to larger datasets with lower quality remains an open question as the computational requirements are vastly out of reach of this project. While ID methods have been successfully scaled to larger datasets [62], data curation is a vital aspect to learn rich representations, which will also be necessary to leverage the full potential of the MIM-Refiner refinement process.
Appendix B Extended Results
B.1 Extended Representation Quality Comparison
We compare linear probing, 1-shot classification performance and runtime against various image models (including non-public ones) in Figure 7.
B.2 Where to Attach ID Heads?
Table 7 ablates different choices of where to attach ID heads on a D2V2 pre-trained ViT-L/16 and ViT-H/14. On ViT-L/16, there is almost no difference of where to attach the ID heads in the last third of blocks. We tried to transfer this insight to ViT-H/14 where the default setting of attaching ID heads to all later blocks performs better. We therefore use the default setting of attaching ID heads to the last 8 blocks (ViT-L and ViT-2B) or the last 12 blocks (ViT-H).
@@ B.2 Where to Attach ID Heads?
B.3 Freezing Early Blocks
Freezing early blocks as a form of regularization to preserve MIM features (similar to related works [50, 44]) is not necessary, as shown in Table 8. Note that we still use a layer-wise learning rate decay [18].
#Frozen $k$-NN
0 81.0
6 80.9
12 80.6
Table 8: Freezing early blocks is not necessary and slightly decreases performance. Ablation conducted with D2V2-L/16. We freeze the first 6 layers to refine MAE-2B to save memory/compute.
B.4 Intermediate Representation Analysis of Additional MIM Methods
(a) MAE [31] $k$-NN classification
(b) MAE [31] segmentation probe
(c) data2vec 2.0 [6] $k$-NN classification
(d) CrossMAE [28] $k$-NN classification
(e) dBOT [55] $k$-NN classification
(f) CAE [15] $k$-NN classification
Figure 8: Intermediate representation analysis of various MIM models. As models get bigger, the $k$-NN accuracy degrades more towards the end of the encoder, especially for L/16 and H/14 models. This pattern is consistent over various MIM models and across tasks. $k$-NN accuracy is calculated on ImageNet-1K [21]. (b) shows the mIoU of linear segmentation probes on ADE20K [97].
B.5 Low-shot Classification Without Color Augmentations
@@
Table 9: ImageNet-1K low-shot and feature evaluations of a D2V2-Refined-H/14 with and without color augmentations. Omitting color augmentations improves ImageNet-1K low-shot performance.
B.6 Extended ImageNet-1K Low-shot and Feature Evaluations
@@
L/16 CrossMAE 16.8 34.0 52.4 63.2 77.7 74.4 53.4
MAE 14.3 34.9 56.9 67.7 79.3 77.5 60.6
dBOT 28.0 46.9 62.4 70.0 80.2 77.8 61.3
D2V2 24.1 58.8 72.1 75.1 81.5 78.2 51.8
CAE 28.1 56.5 68.4 71.4 79.5 80.0 66.9
iBOT 48.5 58.2 66.5 73.3 79.0 81.1 78.0
Mugs 52.9 62.3 69.4 76.2 80.3 82.1 80.4
@@
g/14 DINOv2 (142M) 60.5 68.3 74.4 79.1 83.8 86.5 83.5
B.7 Extended ImageNet-1K Cluster Evaluations
@@
(a) Cluster Accuracy
(b) Silhouette Score
(c) Cluster Similarity
Figure 9: Block-wise cluster analysis for refined and unrefined MAE and D2V2 (H/14). (a): Cluster accuracy w.r.t. lowest $k$-means loss per block. (b): Silhouette score w.r.t. ground truth ImageNet-1K classes. (c): Pairwise cluster label similarity between subsequent blocks (details in Section B.7).
B.8 Multi-model Clustering Evaluations
We evaluate unsupervised cluster accuracy of MIM-Refiner in combination with other foundation models using the TURTLE [29] framework. The results in Table 13 show that MIM-Refiner learns features that are complementary to features learned from foundation models. Combining multiple MIM-Refiner models boosts unsupervised classification accuracy of individual models surpassing the best $k$-means accuracy of D2V2-Ref.-H/14 (67.3%). TURTLE with the feature spaces of MAE-Ref., dBOT-Ref. and D2V2-Ref improves the state-of-the-art of unsupervised classification accuracy using only ImageNet-1K for pre-training to 71.6%. When additionally including foundation models that were pre-trained on web-scale data, MIM-Refiner consistently improves performance. MAE-Refined in combination with DINOv2 [62], CLIP [68] and SWAG [73] sets a new state-of-the-art of 76.8%.
We conduct this study by implementing MIM-Refiner models into the official implementation of TURTLEhttps://github.com/mlbio-epfl/turtle and train with the recommended default settings. We do not tune any hyperparameters.
ImageNet-1K pre-training Web-scale pre-training
MAE-Ref. dBOT-Ref. D2V2-Ref. DINOv2 CLIP SWAG ACC
MIM-Refiner only
61.7
62.2
70.4
71.6
MIM-Refiner + DINOv2
68.5
58.3
74.0
73.7
MIM-Refiner + DINOv2 + CLIP
72.9
69.4
75.0
75.0
MIM-Refiner + DINOv2 + CLIP + SWAG
74.8
74.9
76.4
76.8
Table 13: Unsupervised classification evaluation using the TURTLE [29] framework. MIM-Refiner models synergize well with foundation models such as DINOv2 [62], CLIP [68] and SWAG [73]. MIM-Refiner in combination with TURTLE [29] sets a new state-of-the-art in ImageNet-1K unsupervised classification without additional data (71.6%) and with additional data (76.8%). Model sizes are H/14 for MIM-Refiner and SWAG [73], g/14 for DINOv2 [62] and L/14 for CLIP [68].
B.9 COCO Object Detection and Instance Segmentation
[…]
MAE dBOT D2V2
[…]
MIM 55.2 49.1 54.7 48.7 56.0 49.5
MIM-Refiner 54.9 48.8 54.6 48.6 55.5 49.3
Difference -0.3 -0.3 -0.1 -0.1 -0.5 -0.2
Table 14: Fine-tuning results for L/16 models on COCO object detection and instance segmentation.
B.10 Extended Transfer Learning Results
@@
L/16 CrossMAE 5.4 45.6 63.7 39.8 81.2 30.9
MAE 7.1 51.6 68.9 42.8 83.0 33.6
dBOT 7.1 53.1 71.2 44.6 83.4 34.1
D2V2 5.5 53.1 70.6 39.4 81.5 38.8
CAE 7.4 54.5 71.2 46.4 84.8 36.5
iBOT 15.8 51.5 65.5 56.0 87.6 35.6
Mugs 19.5 53.2 66.9 61.5 87.9 34.8
@@
L/7 MSN 17.0 38.0 48.1 - - -
B.11 VTAB Individual Dataset Results
@@
B.12 Extended Comparison of Fine-tuning on ImageNet-1K and iNat18
Table 18 extends Table 5 with additional MIM models. MIM-Refiner consistently improves also on fine-tuning with an abundance of labels. Table 19 shows individual performances on robustness and domain generalization benchmarks.
@@ B.12 Extended Comparison of Fine-tuning on ImageNet-1K and iNat18
B.13 Fine-tuning ViT-2B Models
@@
B.14 Impact of Multi-Crop Augmentation
@@ B.15 High-dimensional $k$-NN
B.16 MIM-Refiner on Smaller Models
MIM-Refiner builds on pre-trained MIM models which excel at larger scales (ViT-L and upwards), we mainly focus on large-scale models in our paper. However, MIM-Refiner also improves MIM models on smaller scales (ViT-B). Additionally, combinations of MIM and ID methods have been explored in various works [41, 83, 93]. However, as these methods introduce significant runtime overhead over MIM models, they mainly focus on smaller models (ViT-L and smaller) where the pre-trained models are also often not published, which makes a comprehensive comparison against these models impossible. Nevertheless, we show that MIM-Refiner is complementary to these methods by refining a Contrastive MAE [41] ViT-B/16 model with MIM-Refiner. Table 22 shows that MIM-Refiner also significantly improves representation quality on smaller models. For training ViT-B models, we only use a single ID head after the last encoder block.
ViT-B/16 $k$-NN 5-shot 2-shot 1-shot
MAE 51.1 43.1 27.1 14.0
MAE-Refined 76.6 64.5 58.6 50.0
CMAE 76.7 43.3 31.2 21.7
CMAE-Refined 78.5 70.1 65.7 57.9
Table 22: MIM-Refiner also significantly improves ViT-B models. Methods that improve MIM by incorporating ID already during pre-training are orthogonal to MIM-Refiner where the refinement process also significantly improves representation quality of these models.
B.17 Runtime Overhead of ID Queue
We benchmark the overhead of the queue with its topk lookup in Table 23. The queue only adds a small amount of overhead as it is only needed for the forward pass (not for the backward pass) and the queue operates in the bottleneck dimension (256 for all models) of the ID head. Our setup for the queue closely follows the one introduced in  [25]. We estimate the runtime of three configurations where the last one is used for training MIM-Refiner models: (i) no queue (ii) queue with top1 NN-swap and (iii) queue with top20 NN-swap. We train these configurations for a short amount of time on a single GPU with the maximal possible batchsize and report the average runtime per sample. Note that the overhead is so small that random runtime fluctuations that are common in modern GPU setups can slightly distort the results.
Queue size topk L/16 H/14 2B/14
0 - 12.8s 30.1s 76.7s
65K 1 12.9s 30.2s 76.9s
65K 20 13.0s 30.5s 77.5s
Table 23: Runtime per sample for different queue configurations. The queue adds only a small amount of overhead. MIM-Refiner uses a queue size of 65K with a top20 NN lookup.
Appendix C Implementation Details
C.1 Evaluations
@@ C.4 ID Head Architecture
C.5 Refinement Hyperparameters
Hyperparameters for the refinement stage are listed in Table 24. Following MAE-CT [50], we initialize all ID heads first by training them with a frozen encoder to ensure a good learning signal from the start of the refinement process. For this initialization, we use the same hyperparameters as in Table 24 except that we use 20 epochs for all models, a learning rate of 2e-4 and a top1-NN lookup. As we do not use a momentum encoder during training, we instead track an EMA of the encoder and use the EMA then for downstream tasks. As ViT-2B is very expensive to train, we freeze the first 6 blocks (for refinement and also for evaluation). As shown in Table 8 this slightly reduces performance but also reduces memory consumption and runtime.
@@ C.5 Refinement Hyperparameters
   Freeze Blocks 0 (L/H), 6 (2B)
@@
   Local View Resolution 96 (L), 98 (H, 2B)
@@
Table 24: MIM-Refiner hyperparameters.
Appendix D Evaluation Details
D.1 GPU Hours Benchmark
For benchmarking GPU hours, we follow the setup from MAE-CT [50]: we conduct a comparison by implementing all methods in a single code-base and conducting short training runs on a single A100 40GB PCIe card. These runs are executed in mixed precision training mode and with the highest possible batchsize that is a power of 2. The runtime of these benchmark runs is then extrapolated to the reported number of epochs. Benchmarks are conducted in pytorch 2.1 with CUDA 12.1. FLOPS are measured with the fvcore libraryhttps://github.com/facebookresearch/fvcore. For the 1-shot classification plot in Figure 2, we do not take into account that some models train on higher resolutions (e.g. DINOv2) for visual clarity.
D.2 MAE Intermediate Representation Analysis
@@
For the visualization in Figure 3(d), we calculate the delta from one block to the next. We do this for both the $k$-NN accuracy and the reconstruction loss. Additionally, we divide by the maximum delta of each metric to transform both metrics into a similar value range and upper bound 1. Figure 11 shows the reconstruction loss per block and the same plot for a MAE L/16, where a similar behavior can be observed.
We conduct the same analysis for refined models and show results in Figure 12.
(a) Reconstruction Loss
(b) $k$-NN Accuracy
Figure 11: (a): Reconstruction loss per block of MAEs. (b) Relative improvement of reconstruction loss and $k$-NN accuracy for a MAE L/16.
(a) Reconstruction Loss
(b) $k$-NN Accuracy
(c) Relative Improvement
Figure 12: Feature analysis of refined models. (a) The reconstruction loss increases in later blocks after refinement. (b) Representation quality increases drastically in later encoder blocks. (c) The relative improvements of MAE-Refined-H/14 shows that the refinement improves representation quality at the cost of reconstruction loss, which benefits downstream performance.
D.3 ImageNet-1K Low-shot Evaluation Details